OpenStudy (anonymous):

In solving the equation (x + 4)(x – 7) = -18, Eric stated that the solution would be x + 4 = -18 => x = -22 or (x – 7) = -18 => x = -11 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning

OpenStudy (anonymous):

@JakeV8

OpenStudy (lukebluefive):

I wouldn't go about it that way. Instead, expand the left hand side of the equation: \[(x+4)(x-7) = x^2-3x-28\] Now add 18 to both sides of the original equation: \[x^2-3x-10 = 0\] Now factor this to get an equation that is equal to zero. Then, the proper values for x will be clearly visible.

OpenStudy (anonymous):

@LukeBlueFive :) Good thing you wouldn't go about it that way... the "Eric" in the problem was wrong :) You can't solve a problem like his initial problem by considering only one parenthesis at a time to find the x values of the solution. That works, but you have to have "zero" on the right side of the equation... he had "-18"

OpenStudy (anonymous):

@JakeV8 explain lol

OpenStudy (anonymous):

So @jdorta1, the way @LukeBlueFive shows is exactly right. What Eric should have done was first multiply out the stuff on the left, then get the "-18" over on the left also to leave just zero on the right. Then, you can factor and solve for "x".

OpenStudy (anonymous):

@JakeV8 ok so ya..how? lol

OpenStudy (anonymous):

Does that make sense? what I said in words is precisely what @LukeBlueFive showed above.

OpenStudy (lukebluefive):

Do you need help factoring the equation?

OpenStudy (anonymous):

ugh math is crazy!

OpenStudy (anonymous):

wow... lots of lag in the screen updates here :( @jdorta1, how about you let @LukeBlueFive walk you through the steps?

OpenStudy (anonymous):

@JakeV8 ok

OpenStudy (anonymous):

so i now have (x+4)(x-7)=x^2-3x-28

OpenStudy (lukebluefive):

Okay, to factor the equation we have: \[x^2 - 3x - 10\] Since we want to factor this into an equation of the form: \[(x+a)(x-b)\] We need the factors of 10, which are 2 & 5. Exactly one of them must be a negative. Which one should we choose, though? Well, the one where their sum will equal -3, so that we will get our middle term -3x.

OpenStudy (anonymous):

gotta back up..how did you get a -10?

OpenStudy (anonymous):

did i lose you guys?

OpenStudy (anonymous):

(x + 4)(x – 7) = -18 <<<<---- Original Now multiply out the left side: x^2 + 4x -7x - 28 = -18 Simplify left side (combine x terms): x^2 - 3x - 28 = -18 Now add 18 to both sides to get rid of the -18 on the right side: x^2 - 3x - 28 + 18 = -18 + 18 Simplify: x^2 - 3x -10 = 0 Now you can factor :)

OpenStudy (anonymous):

@JakeV8 ok so how do you factor?

OpenStudy (anonymous):

Rather than re-type it, can you review what @LukeBlueFive a few responses ago? that was a good explanation of how to factor.

OpenStudy (anonymous):

@JakeV8 so it would be (x+2)(x-5)

OpenStudy (anonymous):

yes... if you multiply that out, it matches what you had... (x+2)(x-5) = x^2 +2x - 5x - 10 which is just x^2 -3x -10... so that's correctly factored.

OpenStudy (anonymous):

ok so is that the answer then?

OpenStudy (lukebluefive):

Now you just need to find what values of x make (x+2)(x-5) = 0.

OpenStudy (anonymous):

almost :) When you have a factored form on the left and " = 0" on the right, you are close... you just need the solution values for x. (x+2)(x-5) = 0 If either (x+2) or (x-5) is = 0, then the stuff on the left multiplies to 0, making the equation true. (x+2) = 0 when x = -2 and (x-5) = 0 when x = 5 So those are the solutions.

OpenStudy (anonymous):

ooooooooh gotcha!

OpenStudy (anonymous):

yay thank you guys so much!

OpenStudy (anonymous):

So, take one quick look at the original problem...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The reason Eric's approach was wrong was that he did that same thing I just did, but he did it with -18 on the other side of the equals sign.

OpenStudy (anonymous):

It has to be zero on the other side.

OpenStudy (anonymous):

ok

OpenStudy (lukebluefive):

@jdorta1 You're very welcome! Glad I could help. BTW, nice assist @JakeV8!

OpenStudy (anonymous):

so the answer is y=-2 and x=5

OpenStudy (anonymous):

So @LukeBlueFive showed the way to get zero on the right side... multiply everything out on the left, then move the -18 over, then refactor it all. Once it's factored on the left, and all = zero on the right, then you can do what I just showed you to find the solutions for x.

OpenStudy (anonymous):

sweet! so who wants the medal thingy lol you guys were both a great help

OpenStudy (anonymous):

Give it to @LukeBlueFive :)

OpenStudy (lukebluefive):

Jake should get it. I had to bail half-way through.

OpenStudy (anonymous):

ok =)

OpenStudy (anonymous):

too late haha @LukeBlueFive got it

OpenStudy (lukebluefive):

lol, well, thank you! ^_^

OpenStudy (anonymous):

=) very welcome.

OpenStudy (anonymous):

hope to have you guys help me again soon =) @JakeV8 and @LukeBlueFive

OpenStudy (anonymous):

Glad to help... :)

OpenStudy (anonymous):

=)

OpenStudy (lukebluefive):

Hope we're available when you need help. =)